# On the topological ranks of Banach -algebras associated with groups of subexponential growth

### v2

@@ 3 Applications to symmetrized twisted $L^{p}$-crossed products
We remit the reader to [AuOr22, Section 3] or [BaKw25, Sections 2,3] for a careful discussion of the algebra $F^{q}_{*}(G,X,\alpha,\omega)$ and of these properties that we just mentioned. In any case, Theorem 1.2 applies to $F^{q}_{*}(G,X,\alpha,\omega)$ and we immediately obtain Theorem 1.4 as a consequence. The next theorem compiles consequences and explicit computations derived from these theorems.
@@
The following are true.
@@
    Let $G$ be a nilpotent group and let $G^{f}\subset G$ denote the (normal) subgroup of torsion elements. Then, for $p\in[1,2]$,
@@
All the affirmations follow from Theorem 1.2, Corollary 2.11 and the fact that the appropriate $C^{*}$-counterpart is true. Indeed, (i) follows from [Ri83, BP91], (ii) follows from [Ka93, Theorem 2], (iii) follows from [CP05], (iv) follows from [Pu90, BlKuRo92] and (v) follows from [LiLi18, Corollary 1.3]. ∎
@@ Acknowledgments
The author thankfully acknowledges support by the NSF grant DMS-2144739. He is grateful to Professors Ben Hayes and Hannes Thiel for their helpful comments and to Iason Moutzouris, who pointed him to Proposition 2.6. The author also thanks Professors Jianguo Zhang and Bartosz Kwaśniewski, who pointed him to the references [LiNi23] and [BaKw25], respectively.
[…]
ADDRESS
Felipe I. Flores
Department of Mathematics, University of Virginia,
114 Kerchof Hall. 141 Cabell Dr,
Charlottesville, Virginia, United States
E-mail: hmy3tf@virginia.edu

### v3

@@ 3 Applications to symmetrized twisted $L^{p}$-crossed products
The completion $F^{p}(G,X,\alpha,\omega):=\overline{\ell^{1}_{\alpha,\omega}(G,C_{0}(X))}^{\|\cdot\|_{p,{\rm max}}}$ is usually called the full $L^{p}$-crossed product associated with $(G,C_{0}(X),\alpha,\omega)$. It is easily seen to be a Banach algebra [^2] The authors in [5, 13] also consider other, more general, completions of the same kind. That is, completing $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$ with respect to certain classes of representations. However, their general definitions only yield seminorms instead of norms. Here, the special feature of the $L^{p}$-case is the existence of regular representations, which are injective.. In the case where the action is trivial, we set $F^{p}(G,\omega):=F^{p}(G,\{\rm pt\},1,\omega)$. If the cocycle is also trivial, then we set $F^{p}(G):=F^{p}(G,\{\rm pt\},1,1)$. These algebras are often called the twisted group $L^{p}$-operator algebra and the group $L^{p}$-operator algebra, respectively.
These algebras have appeared in full generality in [5] and in [13]. Note, however, that the untwisted crossed product also appears in [19, 11]. Twisted group algebras appear in [21], while [16] deals exclusively with group algebras. For the groupoid case, we recommend [2].
It is observed in [5, Lemma 2.11] that the formulas
$\pi(f)\xi(x,t)=f(x)\xi(x,t),\quad v_{y}\xi(x,t)=\omega(y,y^{-1}t)(x)\xi(\alpha_{y^{-1}}(x),y^{-1}t),$
valid for $f\in C_{0}(X)$, $x,y,t\in G$, and $\xi\in\ell^{p}(X\times G)$, define a covariant representation of $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$ on $\ell^{p}(X\times G)$. Its integrated form satisfies the formula (note the special notation)
$\Lambda_{p}(\Phi)\xi(x,t)=\sum_{y\in G}\Phi(y)(x)\omega(y,y^{-1}t)(x)\xi(\alpha_{y^{-1}}(x),y^{-1}t).$
Using this representation, one may define the reduced twisted crossed product as the completion $\overline{\Lambda_{p}\big(\ell^{1}_{\alpha,\omega}(G,C_{0}(X))\big)}\subset\mathbb{B}(\ell^{p}(X\times G))$. However, in our setting (the groups are amenable), this construction is known to coincide with the full crossed product [5, Theorem 3.5].
In any case, the representation $\Lambda_{p}$ allows us to introduce a ‘symmetrized’ version of the $L^{p}$-crossed product. This construction appears explicitly in [5]; similar constructions also appear in [2, 15, 16, 13], with the only difference being the degree of generality that these authors allow. In any case, consider the norm
$\lVert\Phi\rVert_{q,*}=\max\{\|\Lambda_{p}(\Phi)\|,\|\Lambda_{q}(\Phi)\|\},$
where $p,q\in[1,\infty]$ are Hölder duals. We define $F^{q}_{*}(G,X,\alpha,\omega)$ to be the completion of $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$ under the norm $\lVert\cdot\rVert_{q,*}$. As before, we define the symmetrized group algebra and the symmetrized twisted group algebra by setting $F^{q}_{*}(G,\omega):=F^{q}_{*}(G,\{\rm pt\},1,\omega)$ and $F^{q}_{*}(G):=F^{q}_{*}(G,\{\rm pt\},1,1)$.
The point is that $F^{q}_{*}(G,X,\alpha,\omega)$ is naturally a Banach -algebra that sits between $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$ and $C_{0}(X)\rtimes_{\alpha,\omega}G$, so it fits our setting perfectly. Indeed, the fact that this algebra has a continuous involution arises from the fact that, for all $\Phi\in\ell^{1}_{\alpha,\omega}(G,C_{0}(X)),\xi\in\ell^{p}(X\times G),\eta\in\ell^{q}(X\times G)$, one has
$\langle\Lambda_{p}(\Phi)\xi,\eta\rangle=\langle\xi,\Lambda_{q}(\Phi^{*})\eta\rangle,$
where $\langle\cdot,\cdot\rangle$ denotes the duality pairing between $\ell^{p}(X\times G)$ and $\ell^{q}(X\times G)$ (cf. [2]). On the other hand, the inclusions $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))\subset F^{q}_{*}(G,X,\alpha,\omega)\subset C_{0}(X)\rtimes_{\alpha,\omega}G$ follow from complex interpolation. Indeed, note that $F^{q}_{*}(G,X,\alpha,\omega)$ acts on both $\ell^{p}(X\times G)$ and $\ell^{q}(X\times G)$, so an immediate application of the Riesz-Thorin interpolation theorem yields the existence of $\theta\in[0,1]$ such that
$\lVert\Lambda_{2}(\Phi)\rVert\leq\lVert\Lambda_{p}(\Phi)\rVert^{1-\theta}\lVert\Lambda_{q}(\Phi)\rVert^{\theta}\leq\lVert\Phi\rVert_{q,*}.$
We remit the reader to [2, Section 3] or [5, Sections 2,3] for a careful discussion of the algebra $F^{q}_{*}(G,X,\alpha,\omega)$ and the properties that we have just mentioned. The interpolation argument can be explicitly found in [2] or in [16]. In [5], a different argument is presented. Note that [5, Definition 4.12] provides a more general construction that takes a set $P\subset[1,\infty]$ as input. In [5, Remark 4.13], the authors provide a sufficient condition for this algebra to be a Banach -algebra.
In any case, Theorem 2.8 and Theorem 2.15 apply to $F^{q}_{*}(G,X,\alpha,\omega)$, while Corollary 2.11 applies to $F^{p}(G,X,\alpha,\omega)$. The following theorem is immediately derived as a consequence.
@@
We finish the article with the following theorem, which compiles some consequences and explicit computations derived from these results.
Theorem 3.3.
The following statements are true.
@@
    Let $G$ be a nilpotent group, and let $G^{f}\subset G$ denote the (normal) subgroup of torsion elements. Then, for $p\in[1,2]$,
@@
All the affirmations follow from Theorem 3.2, Corollary 2.11, and the observation that the appropriate result also holds in the corresponding $C^{*}$-case. Indeed, (i) follows from [33, 9]; (ii) follows from [22, Theorem 2]; (iii) follows from [26]; (iv) follows from [31, 7]; and (v) follows from [24, Corollary 1.3]. ∎
@@ Acknowledgments
The author gratefully acknowledges support by the NSF grant DMS-2144739. He is grateful to Professors Ben Hayes, Hannes Thiel and the anonymous referee for their helpful comments. The author also thanks Iason Moutzouris, who pointed him to Proposition 2.6, and Professors Jianguo Zhang and Bartosz Kwaśniewski, for pointing him to the references [23] and [5], respectively.